OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture?

OpenStudy (anonymous):

700(1000) + 1200(R)=1,000,000 700,000 + 1200R = 1,000,000 -700,000 -700,000 1200R=300,000 1200R/1200 = 300,000/1200 R = 250 Robusta Beans